OpenStudy (anonymous):

Simplify the square root 98 times x to the fifth power. a.2 times x squared times the square root of 7 times x b.7 times x squared times the square root of 2 times x c.7 times x squared times the square root of 2 d.2 times x squared times the square root of 7

OpenStudy (campbell_st):

is the question \[\sqrt{98x^5}\] or \[\sqrt{98}x^5\] \[\sqrt{98x^5} = 7x^2\sqrt{2x}\] or \[7\sqrt{2}x^5\]

OpenStudy (anonymous):

sqrt98 x^5 = 7sqrt2 x ^5